OpenStudy (babyd):

how to convert 54 and 5/8% from percent to a decimal

OpenStudy (anonymous):

percent -> decimal divide percent by 100 54/100=0.54 (5/8)%=5/8/100 = 5/800=0.625

OpenStudy (anonymous):

54 and 5/8% 54 and 0.625% 54.625% 0.54625

OpenStudy (anonymous):

to convert 54% to decimal divide by 100 that is .54.. and 5/8%=5/8/100=500/8=250/4=125/2=62.5

OpenStudy (babyd):

thanks

OpenStudy (babyd):

could any of you solve another one r 2 like 7 and 1/2 converting froma fraction to percent

OpenStudy (radar):

7 and 1/2=7.5 to convert to percent move decimal two places to right for 750%